An asset was purchased for $71,000 and originally estimated to have a useful life of 10 years with a residual value of $3,500. After two years of straight-line depreciation, it was determined that the remaining useful life of the asset was only two years with a residual value of $1,400.

a) Determine the amount of the annual depreciation for the first two years.
$

b) Determine the book value at the end of Year 2.
$

c) Determine the depreciation expense for each of the remaining years after revision.
$
Answer #1

A) Annual depreciation = (Cost - Salvage value) /Useful life

= (71,000-3,500)/10

= 6750

B) Book value at the end of year 2

= 71,000 - (6750*2 years)

= 57,500

C) Depreciation expense = (57,500-1400)/2

= 28,050

Comment if you face any issues.